how do i figure out the y intercept and the x intercept in
5x^2 + 5y^2 = 45?

please helpp

Answer by stanbon(75887)

how do i figure out the y intercept and the x intercept in
5x^2 + 5y^2 = 45?
----
Divide thru by 5 to get:
x^2 + y^2 = 9
---
The graph is a circle with center at (0,0)
and radius = sqrt(9) = 3
---
x-intercepts: (-3,0) and (+3,0)
---
y-intercepts: (0,-3) and (0,+3)
==================================
Cheers,
Stan H.